Grade 12Physical Chemistry

Calculate the mole fraction, molarity and molality of NH3 if it is in a solution composed of 30.6 g NH3 in 81.3 g of H2O. The density of the solution is 0.982 g/mL and the density of water is 1.00 g/mL.

ISHAN KAPADI
5 Years agoGrade 12

1 Answer

Dr Makda
5 Years ago
First of all let's calculate the molar weight of 
NH3 =14+ 3= 17g
H2O =2 + 16= 18g 
Now let's calculate the number of moles of 
NH3 in 30.6g by dividing it by molar weight = 30.6/17 = 1.8 moles
Similarly, the number of moles of H20 in 81.3g = 81.3/18 = 4.5
 
Mole fraction = number of moles of solute/number of moles of solvent = 1.8/4.5 = 0.4
Molality = number of moles of solute /weight of solvent in kg = 1.8/81.3 X 1000 = 22 moles per gm of water
Molarity = number of moles of solute / volume of solvent in L = 1.8/81.3 × 1000 = 22 M per mL of water 
Now except for the units, the molarity and molality is the same in this case, this is because, density of water is 1g per mL at room temperature.